OpenStudy (anonymous):

Write the equation of the line in slope-intercept form: m=7/4, (0,3/8)

pooja195 (pooja195):

y=mx+b m=slope b=y-intercept (x,y) if x is 0 then the y value is 3/8 thefore b=3/8

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 0}}\quad ,&{\color{blue}{ \frac{3}{8}}})\quad % (c,d) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{7}{4} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

pooja195 (pooja195):

@jdoe0001 they don't need point slope form.